OpenStudy (anonymous):

I need to factor 2x^3-16

OpenStudy (anonymous):

Factor out and 2 and you will have difference of cube

OpenStudy (anonymous):

2x^3=16 x^3=8 Now try what i tried in the last problem.but this time with 16

OpenStudy (anonymous):

SOMEONE GIVE IT TO ME STEP BY STEP! Haha, thank you. :)

OpenStudy (anonymous):

\[a^3-b^3=(a-b) \left(a^2+a b+b^2\right)\]

OpenStudy (anonymous):

x^3-8 =(x-2)(x^2 +2x-4) = x=2 In general a3 - b3 = (a - b)(a2 + ab + b2)

OpenStudy (anonymous):

So in this problem b=2.52??? and a=x???

OpenStudy (anonymous):

b=2,a=x

OpenStudy (anonymous):

How is b=2????

OpenStudy (anonymous):

you need to try and convert 8 to the form of b^3

OpenStudy (anonymous):

OOOOHHHH! So you divide by 2 first!!! Gotcha now, ok

OpenStudy (anonymous):

Not exactly divide, we try and finding the factors of 8 in such a way they repeat thrice.Like 2*2*2

OpenStudy (anonymous):

I kind of get it......

OpenStudy (anonymous):

Similarly with -27, when you factorise it, you get -3*-3*-3. I mean look for these factors rather than, 3*3*-3

OpenStudy (anonymous):

So when I substitute do I say -8 or 8?

OpenStudy (anonymous):

I get it now, haha

OpenStudy (anonymous):

here, when you frame your equation, it becomes x^3=8. So you try to factoris 8 in terms of a cube not -8

OpenStudy (anonymous):

ok, so 8

OpenStudy (anonymous):

So is my final answer (x+3)(x^2+8x+64)?

OpenStudy (anonymous):

No, b^3 =8 not b. so it will be (x-2)(x^2+2x+4)=0 this when expanded gives x^3-8=0

OpenStudy (anonymous):

OH! So b=2

OpenStudy (anonymous):

yep!